# Well-posedness and stability of the Lagrange representation of the n-D wave equation via boundary triples

### v1

@@ 5. Conclusion
For our results on stability we first point out, that the situation covered in this article does not provide access to the tools usually employed in the verification of stability of solutions, which in most instances means exponential stability. This is because of the simple fact, that in our case, the damping happens on the boundary. Because of that lack of stronger techniques, we cannot show exponential stability and an application of the Gearhart–Prüss theorem seems out of reach, as we would need to prove sufficient resolvent estimates. The notion of “semi-uniform stability” presents a way out, with a convenient criterium (Proposition 4.3) requiring simple spectral theory. From the fact, that the domain of our differential operator is compactly embedded into the state space (Proposition 4.4) we immediately can conclude that its spectrum is a pure point-spectrum, making the spectral condition of Proposition 4.3 relatively easy to check. We point out, that the bottleneck for the stability part is the application of a unique continuation theorem to prove that the invariant system Equation 10 with zero boundary only admits the trivial solution. Only in this instance do we require Lipschitz continuity of $T$ and $\rho$.
Appendix A Unique continuation
In this section we want to briefly explain how we apply the unique continuation principle. In particular we use the version from [10]. To compare the notation we have $A=T$, $\mathbf{b}=0$ and $V=\lambda^{2}\rho$ (left hand side of the equality sign is the notation from [10]).
$\operatorname{div}T\nabla w+\lambda^{2}\rho w$ $=0\quad\text{in}\ \Omega,$
$\nu\cdot T\nabla w$ $=0\quad\text{on}\ \Gamma,$
$w$ $=0\quad\text{on}\ \Gamma.$
Then we define for every $x\in\Omega$ the set $\Omega_{x}$ as in Figure 1, i.e., we choose a smooth path from $x$ to a point outside of $\Omega$ that crosses $\partial\Omega$ in (the interior of) $\Gamma$ and we define $\Omega_{x}$ as a smooth neighborhood of this path that is sufficiently small. Furthermore, we define
$w_{x}(\zeta)\coloneqq\begin{cases}w(\zeta),&\zeta\in\Omega\cap\Omega_{x},\\ 0,&\zeta\in\Omega_{x}\setminus\Omega.\end{cases}$
$x$$\Omega_{x}$$\Omega$$\Gamma$
Note that we can extend $T$ and $\rho$ Lipschitz continuously and boundedly to all of $\mathbb{R}^{d}$. [^15] We can Lipschitz continuously extend $T$ to a neigborhood of $\Omega$ and we then work with $\alpha T+(1-\alpha)\mathrm{I}$, where $\alpha$ is a cutoff function. This construction gives an extension that is globally bounded. Analogously, we can extend $\rho$. By splitting the area of integration, using integration by parts and the boundary conditions on $\Gamma$ we can see that $w_{x}\in\mathrm{H}^{1}(\Omega_{x})$ such that $T\nabla w_{x}\in\mathrm{H}(\operatorname{div},\Omega_{x})$. In particular, this leads to $w_{x}$ solving
We aim to apply the unique continuation principle [10, Thm. 1.1] to conclude that $w_{x}$ is zero. Note that the assumptions of [10, Thm. 1.1] are easy to check for our setting as $\mathbf{b}=0$ and $V=\lambda^{2}\rho$ is bounded. Hence, $w_{x}=0$ and consequently also $w=0$.
Appendix B Regularity
Note that for general (non-Lipschitz continuous) $T$ the $\mathrm{C}^{\infty}$ functions are not necessarily dense in the domain of $\operatorname{div}T\nabla$, as they may not even lie in the domain. Hence, arguments that rely on smooth functions in the domain of $\operatorname{div}T\nabla$ and extend these properties by density require an alternative dense set of regular functions. In a previous version we relied on such arguments, but we could find more direct methods. Nevertheless, we aim to identify such a set of “regular” functions that is dense for potential future applications. In particular we are concerned with boundary regularity—specifically, the existence of a meaningful $\mathrm{L}^{2}(\partial\Omega)$ normal trace for $T\nabla w$. We therefore define the following spaces
$\mathring{\mathrm{H}}(\operatorname{div}0,\Omega)$$\coloneqq\{u\in\mathrm{H}(\operatorname{div},\Omega)\,|\,\mathopen{}\operatorname{div}u=0,\gamma_{\nu}u=0\}$
$\hat{\mathrm{H}}(\operatorname{div},\Omega)$$\coloneqq\{u\in\mathrm{H}(\operatorname{div},\Omega)\,|\,\mathopen{}\gamma_{\nu}u\in\mathrm{L}^{2}(\partial\Omega)\}.$
We will disregard the additional boundary condition on $\Gamma_{0}$ imposed by $\mathring{\mathrm{H}}_{\Gamma_{0}}^{1}(\Omega)$ (i.e., we regard $\Gamma_{0}=\emptyset$) as it actually simplifies the argument rather than complicating it, because in this space $\lVert\nabla\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$ is equivalent to the full $\mathrm{H}^{1}(\Omega)$ norm. The only part that needs more attention in that case is the Helmholtz decomposition, but this is covered by [3, Thm. 5.3].
Theorem B.1.
The set $D_{1}\coloneqq\{f\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}T\nabla f\in\hat{\mathrm{H}}(\operatorname{div},\Omega)\}$ is a core of $L=\operatorname{div}T\nabla$ (with $\operatorname{dom}L=\{f\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}T\nabla f\in\mathrm{H}(\operatorname{div},\Omega)\}$ as operator on $\mathrm{L}^{2}(\Omega)$).
Proof.
Let $f\in\operatorname{dom}L$. Note that $\hat{\mathrm{H}}(\operatorname{div},\Omega)$ is dense in $\mathrm{H}(\operatorname{div},\Omega)$ w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}(\operatorname{div},\Omega)}$. Hence, for given $\epsilon>0$ there exists a $g\in\hat{\mathrm{H}}(\operatorname{div},\Omega)$ such that $\lVert T\nabla f-g\rVert_{\mathrm{H}(\operatorname{div},\Omega)}\leq\epsilon$. Moreover, similar to Theorem 2.6, by replacing $\mathring{\nabla}$, the gradient with Dirichlet boundary, with $\nabla$, as in the proof there we can decompose $\mathrm{L}^{2}(\Omega)$ into
where $a\perp_{T^{-1}}b$ means $\langle T^{-1}a,b\rangle_{\mathrm{L}^{2}(\Omega)}=0$. Hence, there exist $h\in\mathrm{H}^{1}(\Omega)$ and $k\in\mathring{\mathrm{H}}(\operatorname{div}0,\Omega)$ such that $g=T\nabla h+k$. Since $g,k\in\hat{\mathrm{H}}(\operatorname{div},\Omega)$ we conclude that also $T\nabla h\in\hat{\mathrm{H}}(\operatorname{div},\Omega)$. Moreover,
and
This gives $\lVert\operatorname{div}T\nabla(f-h)\rVert_{\mathrm{L}^{2}(\Omega)}=\lVert\operatorname{div}T\nabla f-g\rVert_{\mathrm{L}^{2}(\Omega)}\leq\epsilon$. Clearly, we also have $\lVert T\nabla f-g\rVert_{\mathrm{L}^{2}(\Omega)}\leq\epsilon$. Moreover, by orthogonality we obtain
$\epsilon^{2}$ $\geq\lVert T\nabla f-T\nabla h-k\rVert_{\mathrm{L}^{2}(\Omega)}^{2}\geq\frac{1}{\lVert T^{\frac{1}{2}}\rVert^{2}}\lVert T\nabla(f-h)-k\rVert_{T^{-1}}^{2}$
$=\frac{1}{\lVert T^{\frac{1}{2}}\rVert^{2}}\bigl{(}\lVert T\nabla(f-h)\rVert_{T^{-1}}^{2}+\lVert k\rVert_{T^{-1}}^{2}\bigr{)},$
Moreover, we have the decomposition
$\mathrm{H}^{1}(\Omega)=\mathbb{C}^{\perp}\oplus\mathbb{C}$
and that $\nabla\colon\mathbb{C}^{\perp}\to\operatorname{ran}\nabla$ is a boundedly invertible mapping, where $\operatorname{ran}\nabla$ is endowed with $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$. Note that $f$ and $h$ can be decomposed according to this decomposition into $f=f_{1}+f_{2}$ and $h=h_{1}+h_{2}$, where $f_{1},h_{1}\in\mathbb{C}^{\perp}$ and $f_{2},h_{2}\in\mathbb{C}$. Since $T$ is also boundedly invertible we obtain
Hence, we define $\phi=h_{1}+f_{2}\in\mathrm{H}^{1}(\Omega)$, which implies $\nabla\phi=\nabla h$ and in particular $\phi\in D_{1}$. This gives
for a constant $C>0$. Note that by construction $\phi\in D_{1}$ and $\lVert f-\phi\rVert_{\operatorname{dom}L}\leq\hat{C}\epsilon$, which proves that $D_{1}$ is a core of $L$. ∎
Consequently for $A=\begin{pmatrix}0&\tfrac{1}{\rho}\\ \operatorname{div}T\nabla&0\end{pmatrix}$ with $\Gamma_{0}=\emptyset$ from Section 3 we obtain:
Corollary B.2.
The set
$D\coloneqq\left\{\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\in\operatorname{dom}(A)\,\middle|\,\mathopen{}\gamma_{\nu}T\nabla x_{1}\in\mathrm{L}^{2}(\partial\Omega)\right\}=D_{1}\times\rho\mathrm{H}^{1}(\Omega)$
is a core of $A$.
Remark B.3.
Note the general statement with non-empty $\Gamma_{0}$ can be obtained relatively similar. The only difference is that in the proof of Theorem B.1 we need the corresponding Helmholtz decomposition
see, e.g., [3, Thm. 5.3].
Appendix C Friedrichs/Poincaré inequality
We use a slightly modified version of Friedrichs/Poincaré’s inequality.
Theorem C.1.
@@ Appendix C Friedrichs/Poincaré inequality
Let $\Omega$ be a bounded and connected Lipschitz domain, $\Gamma\subseteq\partial\Omega$ with positive measure and $k_{1}\in\mathcal{L}_{\mathrm{b}}(\mathrm{L}^{2}(\Gamma))$ such that $1\notin\ker k_{1}$. Then there exists a $C>0$ such that for all $f\in\mathrm{H}^{1}(\Omega)$
@@
We define $g_{n}\coloneqq\frac{f_{n}}{\lVert f_{n}\rVert_{\mathrm{L}^{2}(\Omega)}}$, which implies
@@
By the Rellich–Kondrachov theorem and passing over to a subsequence the sequence $(g_{n})_{n\in\mathbb{N}}$ converges to a $g\in\mathrm{L}^{2}(\Omega)$ (w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$). For $\phi\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega)$ we have
@@
which implies that $g\in\mathrm{H}^{1}(\Omega)$ and $\nabla g=0$. Consequently, $g$ is constant, i.e., there exists a $c\in\mathbb{C}$ such that $g=c$. Moreover, the sequence $(g_{n})_{n\in\mathbb{N}}$ converges also w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$ to $g$. This gives
$k_{1}c=k_{1}\gamma_{0}g=\lim_{n\to\infty}k_{1}^{\frac{1}{2}}k_{1}^{\frac{1}{2}}\gamma_{0}g_{n}=0$
@@
and from the assumption on $k_{1}$ we conclude $c=0$, which contradicts $\lVert g\rVert_{\mathrm{L}^{2}(\Omega)}=1$. ∎

### v2

@@ 5. Conclusion
For our results on stability we first point out, that the situation covered in this article does not grant access to the tools usually employed in the verification of stability of solutions, which in most instances means exponential stability. The simple reason is, that in the considered case here, the damping happens on the boundary. Because of the lack of stronger techniques, we cannot show exponential stability and an application of the Gearhart–Prüss theorem seems out of reach, as we would need to prove sufficient resolvent estimates. The notion of “semi-uniform stability” presents a way out, with a convenient criterium (Proposition 4.3) requiring simple spectral theory. From the fact, that the domain of our differential operator $A_{\Theta}$ is compactly embedded into the state space (Proposition 4.4), we can immediately conclude that its spectrum is a pure point-spectrum, making the spectral condition of Proposition 4.3 easy to check. We point out, that the bottleneck for the stability part is the application of a unique continuation theorem to prove that the invariant system (11) with zero boundary only admits the trivial solution. Only in this instance do we require Lipschitz continuity of $T$ and $\rho$.
Appendix A Friedrichs/Poincaré inequality
We use a slightly modified version of Friedrichs/Poincaré inequality.
Theorem A.1.
@@ Appendix A Friedrichs/Poincaré inequality
Let $\Omega$ be a bounded and connected Lipschitz domain, $\Gamma\subseteq\partial\Omega$ of positive measure and $k_{1}\in\mathcal{L}_{\mathrm{b}}(\mathrm{L}^{2}(\Gamma))$ such that $1\notin\ker k_{1}$. Then there exists a $C>0$ such that for all $f\in\mathrm{H}^{1}(\Omega)$:
@@
We define $g_{n}\coloneq\frac{f_{n}}{\lVert f_{n}\rVert_{\mathrm{L}^{2}(\Omega)}}$, which implies
@@
By the Rellich–Kondrachov theorem ([19, Thm. 12.30]) and after passing to a subsequence, $(g_{n})_{n\in\mathbb{N}}$ converges to a $g\in\mathrm{L}^{2}(\Omega)$ (w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$). For $\phi\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega)^{d}$ we have
@@
which implies that $g\in\mathrm{H}^{1}(\Omega)$ and $\nabla g=0$. Consequently, $g$ is constant, i.e., there exists a $c\in\mathbb{C}$ such that $g=c$. Moreover, the sequence $(g_{n})_{n\in\mathbb{N}}$ converges also w.r.t. $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}$ to $g$. This shows
$k_{1}c=k_{1}\gamma_{0}g=k_{1}^{\frac{1}{2}}\lim_{n\to\infty}k_{1}^{\frac{1}{2}}\gamma_{0}g_{n}=0$
@@
and from the assumption on $k_{1}$ we deduce $c=0$, which contradicts $\lVert g\rVert_{\mathrm{L}^{2}(\Omega)}=1$. ∎
Appendix B Regularity
Note that for general (non-Lipschitz continuous) $T$, the $\mathrm{C}^{\infty}$-functions are not necessarily dense in the domain of $\operatorname{div}T\nabla$, as they may not even lie in the domain. Hence, arguments that rely on smooth functions in the domain of $\operatorname{div}T\nabla$ and extend these properties by density, require an alternative dense set of regular functions. It is possible to obtain (some of) the results of Section 3 by identifying such a set of functions, but ultimately it is not required for our purposes. Nevertheless, we present a statement identifying a core for $A$ that can be viewed as a regularity statement for the boundary; specifically, about the existence of a meaningful $\mathrm{L}^{2}(\partial\Omega)$ normal trace for $T\nabla w$. We define the following spaces
$\mathring{\mathrm{H}}(\operatorname{div}0,\Omega)$ $\coloneq\{u\in\mathrm{H}(\operatorname{div},\Omega)\,|\,\mathopen{}\operatorname{div}u=0,\gamma_{\nu}u=0\},$
$\hat{\mathrm{H}}(\operatorname{div},\Omega)$ $\coloneq\{u\in\mathrm{H}(\operatorname{div},\Omega)\,|\,\mathopen{}\gamma_{\nu}u\in\mathrm{L}^{2}(\partial\Omega)\}.$
We will disregard the additional boundary condition on $\Gamma_{0}$ imposed by $\mathring{\mathrm{H}}_{\Gamma_{0}}^{1}(\Omega)$ (i.e., we regard $\Gamma_{0}=\emptyset$), as it actually simplifies the argument rather than complicating it, because in this space, $\lVert\nabla\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$ is equivalent to the full $\mathrm{H}^{1}(\Omega)$-norm. The only non-obvious aspect in treating such a split boundary is to establish a Helmholtz decomposition, but fortunately this situation is covered by [5, Thm. 5.3].
Theorem B.1.
The set $D_{1}\coloneq\{f\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}T\nabla f\in\hat{\mathrm{H}}(\operatorname{div},\Omega)\}$ is a core of $L=\operatorname{div}T\nabla$ (with $\operatorname{dom}L=\{f\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}T\nabla f\in\mathrm{H}(\operatorname{div},\Omega)\}$ as an operator on $\mathrm{L}^{2}(\Omega)$).
Proof.
[…]
where $a\perp_{T^{-1}}b$ means $\langle T^{-1}a,b\rangle_{\mathrm{L}^{2}(\Omega)}=0$. Hence, there exist $h\in\mathrm{H}^{1}(\Omega)$ and $k\in\mathring{\mathrm{H}}(\operatorname{div}0,\Omega)$ such that $g=T\nabla h+k$. Since $g,k\in\hat{\mathrm{H}}(\operatorname{div},\Omega)$, we conclude that $T\nabla h\in\hat{\mathrm{H}}(\operatorname{div},\Omega)$. Moreover,
and
This shows $\lVert\operatorname{div}T\nabla(f-h)\rVert_{\mathrm{L}^{2}(\Omega)}=\lVert\operatorname{div}(T\nabla f-g)\rVert_{\mathrm{L}^{2}(\Omega)}\leq\epsilon$. Clearly, we also have $\lVert T\nabla f-g\rVert_{\mathrm{L}^{2}(\Omega)}\leq\epsilon$. Moreover, by orthogonality we obtain
$\epsilon^{2}$ $\geq\lVert T\nabla f-T\nabla h-k\rVert_{\mathrm{L}^{2}(\Omega)}^{2}\geq\frac{1}{\lVert T^{-\frac{1}{2}}\rVert^{2}}\lVert T\nabla(f-h)-k\rVert_{T^{-1}}^{2}$
$=\frac{1}{\lVert T^{-\frac{1}{2}}\rVert^{2}}\bigl(\lVert T\nabla(f-h)\rVert_{T^{-1}}^{2}+\lVert k\rVert_{T^{-1}}^{2}\bigr),$
Moreover, identifying $\operatorname{ker}\nabla=\mathbb{C}$, we have the decomposition
$\mathrm{H}^{1}(\Omega)=\mathbb{C}^{\perp}\oplus\mathbb{C}$
and that $\nabla\colon\mathbb{C}^{\perp}\to\operatorname{ran}\nabla$ is a boundedly invertible mapping, where $\operatorname{ran}\nabla$ is endowed with $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}$. Note that $f$ and $h$ can be decomposed according to this decomposition into $f=f_{1}+f_{2}$ and $h=h_{1}+h_{2}$, where $f_{1},h_{1}\in\mathbb{C}^{\perp}$ and $f_{2},h_{2}\in\mathbb{C}$. Since $T$ is boundedly invertible, we obtain
We define $\phi=h_{1}+f_{2}\in\mathrm{H}^{1}(\Omega)$ and obtain $\nabla\phi=\nabla h$; in particular $\phi\in D_{1}$. This gives
for a constant $\tilde{C}>0$. Note that by construction, $\phi\in D_{1}$ and $\lVert f-\phi\rVert_{\operatorname{dom}L}\leq\hat{C}\epsilon$, which proves that $D_{1}$ is a core of $L$. ∎
Consequently, for $A=\begin{pmatrix}0&\tfrac{1}{\rho}\\ \operatorname{div}T\nabla&0\end{pmatrix}$ with $\Gamma_{0}=\emptyset$ from Section 3, we obtain:
Corollary B.2.
The set
$D\coloneq\left\{\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\in\operatorname{dom}(A)\,\middle|\,\mathopen{}\gamma_{\nu}T\nabla x_{1}\in\mathrm{L}^{2}(\partial\Omega)\right\}=D_{1}\times\rho\mathrm{H}^{1}(\Omega)$
is a core of $A$.
Remark B.3.
Note that the general statement with non-empty $\Gamma_{0}$ can be obtained relatively similar. The only difference is that in the proof of Theorem B.1 we need the corresponding Helmholtz decomposition
see, e.g., [5, Thm. 5.3].
Appendix C Unique continuation
In this section we want to briefly explain, how we apply the unique continuation principle. In particular, we use the version from [13]. To compare the notation, we have $A=T$, $\mathbf{b}=0$ and $V=\lambda^{2}\rho$ (on the left-hand side of the equality sign is the notation from [13]). We impose the same assumption on $T$ and $\rho$ as in Section 4, i.e., $T$ and $\rho$ are Lipschitz-continuous.
$\begin{aligned} \operatorname{div}T\nabla w+\lambda^{2}\rho w&=0\quad\text{in}\ \Omega,\\ \nu\cdot T\nabla w&=0\quad\text{on}\ \Gamma,\\ w&=0\quad\text{on}\ \Gamma.\end{aligned}$ (12)
We define for every $x\in\Omega$ the set $\Omega_{x}$ as in Figure 1, i.e., we choose a smooth path from $x$ to a point outside of $\Omega$ that crosses $\partial\Omega$ in (the interior of) $\Gamma$ and we define $\Omega_{x}$ as a smooth neighborhood of this path that is sufficiently small (it does not intersect $\partial\Omega$ outside of $\Gamma$). Furthermore, we define
$w_{x}(\zeta)\coloneq\begin{cases}w(\zeta),&\zeta\in\Omega\cap\Omega_{x},\\ 0,&\zeta\in\Omega_{x}\setminus\Omega.\end{cases}$ (13)
$x\Omega_{x}$$\Omega$$\Gamma$
Note that we can extend $T$ and $\rho$ Lipschitz-continuously and boundedly to all of $\mathbb{R}^{d}$. [^16] We can Lipschitz-continuously extend $T$ to a neigborhood of $\Omega$ and we then work with $\alpha T+(1-\alpha)\mathrm{I}$, where $\alpha$ is a cutoff function. This construction gives an extension that is globally bounded. Analogously, we can extend $\rho$. By splitting the area of integration, using integration by parts and the boundary conditions on $\Gamma$, we can see that $w_{x}\in\mathrm{H}^{1}(\Omega_{x})$ such that $T\nabla w_{x}\in\mathrm{H}(\operatorname{div},\Omega_{x})$. In particular, this leads to $w_{x}$ solving (shown in Lemma C.1)
We aim to apply the unique continuation principle [13, Thm. 1.1] to conclude that $w_{x}$ is zero. The assumptions of [13, Thm. 1.1] are easy to check for our setting, since $\mathbf{b}=0$ and $V=\lambda^{2}\rho$ is bounded. Hence, $w_{x}=0$ for every $x\in\Omega$ and consequently also $w=0$.
Lemma C.1.
Let $w\in\mathring{\mathrm{H}}^{1}_{\Gamma}(\Omega)$ be a weak solution of (12) and $w_{x}$ be defined as in (13). Then $w_{x}\in\mathrm{H}^{1}(\Omega_{x})$ and $T\nabla w_{x}\in\mathrm{H}(\operatorname{div},\Omega_{x})$ with
$\nabla w_{x}=\begin{cases}\nabla w(\zeta),&\zeta\in\Omega\cap\Omega_{x},\\ 0,&\zeta\in\Omega_{x}\setminus\Omega,\end{cases}\quad\text{and}\quad\operatorname{div}T\nabla w_{x}=\begin{cases}\operatorname{div}T\nabla w(\zeta),&\zeta\in\Omega\cap\Omega_{x},\\ 0,&\zeta\in\Omega_{x}\setminus\Omega.\end{cases}$
Moreover, $w_{x}$ satisfies $\operatorname{div}T\nabla w_{x}+\lambda^{2}\rho w_{x}=0$ in $\Omega_{x}$.
Proof.
Let $\phi\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega_{x};\mathbb{C}^{d})$. Then $\phi\in\mathrm{C}_{\mathrm{c}}^{\infty}(\mathbb{R}^{d};\mathbb{C}^{d})$ such that $\operatorname{supp}\phi\cap(\partial\Omega\setminus\Gamma)=\emptyset$. This allows us to use integration by parts on $\Omega$ for $w\in\mathring{\mathrm{H}}^{1}_{\Gamma}(\Omega)$ and $\phi$ without boundary terms (weak characterization of $\mathring{\mathrm{H}}^{1}_{\Gamma}(\Omega)$, cf. [5]). Moreover, $\phi$—and therefore also $\operatorname{div}\phi$—is 0 on $\Omega\setminus\Omega_{x}$. In the following, we will only specify the domain of the $\mathrm{L}^{2}$ space in the index of the inner product, e.g., $\langle\boldsymbol{\cdot},\boldsymbol{\cdot}\rangle_{\Omega}$ instead of $\langle\boldsymbol{\cdot},\boldsymbol{\cdot}\rangle_{\mathrm{L}^{2}(\Omega)}$.
$\langle w_{x},\operatorname{div}\phi\rangle_{\Omega_{x}}$ $=\langle\underbrace{w_{x}}_{=\mathrlap{w}},\operatorname{div}\phi\rangle_{\Omega_{x}\cap\Omega}+\langle\underbrace{w_{x}}_{=\mathrlap{0}},\operatorname{div}\phi\rangle_{\Omega_{x}\setminus\Omega}=\langle w,\operatorname{div}\phi\rangle_{\Omega}+\langle w,\underbrace{\operatorname{div}\phi}_{=\mathrlap{0}}\rangle_{\Omega\setminus\Omega_{x}}$
$\stackrel{{\scriptstyle\mathllap{\text{Int.\ by parts}}}}{{=}}-\langle\nabla w,\phi\rangle_{\Omega}=-\langle\nabla w,\phi\rangle_{\Omega_{x}\cap\Omega}-\langle 0,\phi\rangle_{\Omega_{x}\setminus\Omega}.$
Hence, $w_{x}\in\mathrm{H}^{1}(\Omega_{x})$ satisfies
$\nabla w_{x}=\begin{cases}\nabla w(\zeta),&\zeta\in\Omega\cap\Omega_{x},\\ 0,&\zeta\in\Omega_{x}\setminus\Omega.\end{cases}$
We proceed analogously to show that $T\nabla w_{x}\in\mathrm{H}(\operatorname{div},\Omega)$: We note that since $w$ is a weak solution, we automatically get $T\nabla w\in\mathring{\mathrm{H}}_{\Gamma}(\operatorname{div},\Omega)$. [^17] Roughly speaking, $\mathring{\mathrm{H}}_{\Gamma}(\operatorname{div},\Omega)$ is $\{f\in\mathrm{H}(\operatorname{div},\Omega)\,|\,\mathopen{}\gamma_{\nu}f\big|_{\Gamma}=0\}$. The problem is, that the restriction to $\Gamma$ in general is not defined. However, there are two approaches that make this statement rigorous (a “strong” one via completion of regular functions and a “weak” one via inner products and test functions). Both approaches fortunately coincide, cf. [5]. There, the authors use the notation $\mathring{\mathsf{D}}_{\Gamma}(\Omega)$ (and $\mathring{\mathcal{D}}_{\Gamma}(\Omega)$) for this set. Let $\psi\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega_{x};\mathbb{C})$. Then $\psi\in\mathrm{C}_{\mathrm{c}}^{\infty}(\mathbb{R}^{d};\mathbb{C})$ such that $\operatorname{supp}\psi\cap(\partial\Omega\setminus\Gamma)=\emptyset$. Again, this allows us to use integration by parts on $\Omega$ without boundary terms (weak characterization of $\mathring{\mathrm{H}}_{\Gamma}(\operatorname{div},\Omega)$, cf. [5]). Hence,
$\langle T\nabla w_{x},\nabla\psi\rangle_{\Omega_{x}}=\langle T\nabla w,\nabla\psi\rangle_{\Omega}=-\langle\operatorname{div}T\nabla w,\psi\rangle_{\Omega}\\ =-\langle\operatorname{div}T\nabla w,\psi\rangle_{\Omega_{x}\cap\Omega}-\langle 0,\psi\rangle_{\Omega_{x}\setminus\Omega},$
which shows $T\nabla w_{x}\in\mathrm{H}(\operatorname{div},\Omega_{x})$ and
$\operatorname{div}T\nabla w_{x}=\begin{cases}\operatorname{div}T\nabla w(\zeta),&\zeta\in\Omega\cap\Omega_{x},\\ 0,&\zeta\in\Omega_{x}\setminus\Omega.\end{cases}$
Since $\operatorname{div}T\nabla w+\lambda^{2}\rho w=0$ holds, we immediately infer the same identity for $w_{x}$. ∎